OpenStudy (anonymous):

solve for u u+1/u+7 +1 = u+2/u+1

OpenStudy (anonymous):

first multiply by u+1 to get (u+1)^2/(u+7)+u+1=u+2

OpenStudy (anonymous):

now multiply by u+7 (u+1)^2+(u+1)(u+7)=(u+2)(u+7)

OpenStudy (anonymous):

now expand the brackets and but everything to one side and solve it

OpenStudy (anonymous):

would the answer have a u in it or just numbers ?

OpenStudy (anonymous):

just number(s), and there might be 2 answers

OpenStudy (anonymous):

because i get u+3/2 idk how :-/

OpenStudy (anonymous):

ok expanding gives u^2+2u+1+u^2+u+7u+7=u^2+2u+7u+14

OpenStudy (anonymous):

now putting everything to the left u^2+u-6=0

OpenStudy (anonymous):

do you know how to solve this?

OpenStudy (anonymous):

wait let me try it give me a sec

OpenStudy (anonymous):

its all to the left so wouldn't it just stay like that or would it turn to u^3-6=0

OpenStudy (anonymous):

I can see 10 u on the LHS and 9u on the RHS